### v1

@@ Abstract
Resource Prediction, Workflow Management, Computational Efficiency, Scout Jobs, PanDA System, Resource Allocation, High-Performance Computing (HPC), Scientific Workflows, Distributed Computing, Multi-Class Classification, Data-Driven Resource Management
@@ 4.1 Data, Features, Target Variables, and Preprocessing
4.2 Model Architecture and Training
The predictive framework employs a hybrid neural network architecture designed to handle heterogeneous workflow metadata. Categorical features (e.g., PROCESSINGTYPE, FRAMEWORK) are encoded using adaptive embedding layers, where the embedding dimension scales logarithmically with feature cardinality,
$\text{embed\_dim}=\min\left(32,\left\lfloor\log_{2}(v)\right\rfloor+1\right),$
where $v$ denotes the number of unique values in the feature (formula designed for this study) Numerical features are directly incorporated into the model. All processed inputs are concatenated and passed through a stack of dense layers (256, 128, and 64 units) with L2 regularization ($\lambda=10^{-4}$), batch normalization [8, 9], and progressive dropout (40%, 30%, 30%).
For training and evaluation, the dataset is split into training/validation (85%) and test (15%) partitions using stratified sampling to preserve the class distribution of the target variable [10]. The model is trained on mini-batches of 256 samples, with class weights applied to address class imbalance in the target labels [11]. Early stopping is employed, monitoring validation accuracy with a patience of four epochs, and the best model weights are restored to prevent overfitting. Training is terminated immediately if any numerical instability (e.g., NaN loss) is detected. Optimization is performed using the Adam optimizer with a learning rate of $5\times 10^{-5}$ [12]. For multi-class classification tasks, sparse categorical cross-entropy loss is used, while binary cross-entropy is applied for binary targets. This configuration ensures stable convergence and robust generalization across diverse workflow types.
@@ 4.3 Evaluation Metrics and Model Performance
To comprehensively assess model performance, we report a combination of standard metrics suited to both binary and multiclass classification problems. Overall accuracy quantifies the proportion of correctly classified instances and provides a basic indicator of prediction reliability. However, because our resource classification tasks involve substantial class imbalance, we present additional metrics that offer a more granular evaluation [13, 14, 15].
@@
Model Classes Accuracy Macro F1 ROC-AUC PR-AUC
Model 1 4 0.88 0.73 0.98 0.96
Model 2 5 0.86 0.73 0.99 0.95
Model 3 2 0.94 0.94 0.99 0.98
Model 4 5 0.91 0.80 0.99 0.97
Micro-average AUC for multiclass models
Table 3: Summary of Model Performance Across Resource Prediction Tasks
To further evaluate discriminative capability, we use the area under the receiver operating characteristic curve (ROC-AUC), with micro-averaged values provided in Table 3. In the multiclass context, this gives a global view of the model’s ability to separate all classes across threshold settings. Figure 5 shows the ROC curves for all models, visually demonstrating the quality of classification boundaries throughout the resource estimation pipeline. For binary classification tasks, including the I/O intensity model, the ROC curve provides insight into the model’s performance independent of class distribution [15]. Table 3 summarizes the overall quantitative outcomes.
Model 1 (RAM classification) attains 88% accuracy and macro F1 of 0.73, with a high ROC-AUC of 0.98. Notably, classwise results in Table 2 indicate the model achieves especially strong recall in bin1 (0.95) and bin4 (0.92). Model 2 (CPU time classification) matches this overall performance, with 86% accuracy, macro F1 of 0.73, and ROC-AUC of 0.99; the lowest per-class F1 scores (bin3: 0.52, bin4: 0.51) are due to class rarity but remain operationally useful. Model 3, a binary classifier for I/O intensity, obtains 94% accuracy and macro F1 of 0.94, along with a near-perfect ROC-AUC of 0.99. Both “low” and “high” I/O intensity classes are identified robustly, as confirmed by the ROC curve in Figure 5c. For Model 4 (walltime classification), accuracy reaches 91% with macro F1 of 0.80 and ROC-AUC of 0.99, indicating uniformly high performance across the five prediction bins.
@@ 5 System Integration and Workflow Impact
We have developed a custom service-oriented architecture for integrating machine learning–driven resource prediction into the PanDA workload management system. The architecture is currently being evaluated in a dedicated test environment that closely replicates production conditions, enabling thorough validation with realistic scientific workloads. As shown in Figure 6, machine learning components are exposed as microservices within the workflow, deployed in containers to ensure consistent and reproducible execution. This approach facilitates automated resource prediction, precise scheduling, and streamlined workflow progression, with results from the test environment informing future production deployment.
@@ 5.1 Integrated Pipeline Evaluation for PanDA Brokerage
To quantify the effectiveness of our integrated solution, the PanDA brokerage system employs four predictive models that operate concurrently, providing real-time resource allocation estimates for incoming tasks. Assessing the overall system performance requires joint evaluation metrics that capture the collective contributions and interactions of these models within the broader workflow. Such an approach enables a holistic understanding of how the predictive ensemble drives improvements in scheduling accuracy, resource utilization, and workflow efficiency [13]. This integrated evaluation informs future refinements and will guide deployment in the full production environment.
@@ 5.2 Performance Benchmark and Workflow Impact
The integration of machine learning into the PanDA workflow system delivers transformative performance improvements. With sub-second inference latency, the ML pipeline achieves an orders-of-magnitude speedup over the traditional Scout-based approach, which required hours of task execution to estimate resource needs.
Table 5 summarizes the key performance metrics:

### v2

@@ Abstract
Resource Prediction, Workflow Management, Computational Efficiency, Scout Jobs, PanDA System, Resource Allocation, High-Performance Computing (HPC), Scientific Workflows, Distributed Computing, Deep Learning, Multi-Class Classification, Data-Driven Resource Management
@@ 4.1 Data, Features, Target Variables, and Preprocessing
4.2 Model Architectures: Empirical Baseline and Deep Learning
A robust baseline for resource requirement classification was established using gradient-boosted decision trees implemented through the XGBoost framework [8]. Categorical workflow metadata features were label-encoded, and hyperparameters such as maximum tree depth, learning rate, and regularization parameters were tuned via grid search with cross-validation. This well-established approach provides an empirical benchmark due to XGBoost’s proven efficiency and scalability in structured data applications.
Building upon this baseline, a deep learning model was designed to capture complex, nonlinear dependencies across heterogeneous workflow metadata. The model employs adaptive embedding layers for categorical variables, which learn dense vector representations that compactly encode latent feature relationships beyond what fixed encodings provide. The embedding dimension for a categorical feature is defined as
$\text{embed\_dim}=\min(32,\lfloor\log_{2}(v)\rfloor+1),$
where $v$ denotes the feature’s cardinality. Numerical features are concatenated directly with the embedded categorical representations, forming a unified input vector processed through a sequence of fully connected layers. Batch normalization [9] is applied after each layer to stabilize training, while dropout and $L_{2}$ regularization are employed to control overfitting. Final architectural parameters and optimization details are described in Section 4.3.
Due to the computational expense of neural network training, hyperparameter exploration used a random search approach to efficiently traverse the high-dimensional configuration space [10]. The resulting architecture serves as the foundation of the proposed predictive framework for resource requirement estimation in large-scale scientific workflows.
4.3 Model Training and Hyperparameter Tuning
The dataset comprised approximately 3.4M tasks, representing 85% of the initial data. This dataset was further partitioned into training/validation (85%) and test (15%) subsets using stratified sampling to preserve the class distribution across splits [11]. Hyperparameter optimization was performed using three-fold cross-validation combined with a randomized search strategy [10]. The search identified an optimal network configuration consisting of three fully connected layers containing 256, 128, and 64 units, respectively, with progressive dropout rates set to 40%, 30%, and 30%. An $L_{2}$ regularization coefficient of $\lambda=10^{-4}$ was applied to all dense layers’ kernels, and the learning rate was fixed at $5\times 10^{-5}$. Using these hyperparameters, the final model was retrained on the entire 3.4M tasks and evaluated on 0.6M unseen data.
To account for class imbalance in workflow classifications, class weights inversely proportional to class frequencies were incorporated into the loss function [12], ensuring balanced learning across categories. Early stopping with a patience of four epochs based on validation accuracy was applied, and the best-performing weights were restored to prevent overfitting [9].
Training utilized the Adam optimizer [13] due to its adaptive learning rate properties and convergence stability. Sparse categorical cross-entropy was used for multi-class resource classification, and binary cross-entropy for binary I/O predictions.
This combination of stratified sampling, randomized search tuning, class weighting, early stopping, and adaptive optimization ensures both high predictive accuracy and strong generalization capability. Benchmarking against the XGBoost baseline was conducted using identical data partitions for a fair comparative analysis, as presented in Section 4.4.
@@ 4.4 Evaluation Metrics and Model Performance
To comprehensively assess model performance, a combination of standard metrics suitable for both binary and multiclass classification problems was employed. Overall accuracy quantifies the proportion of correctly classified instances and provides a basic indicator of prediction reliability. However, since the resource classification tasks exhibit substantial class imbalance, we present additional metrics that offer a more granular evaluation [14, 15, 16].
Both the deep learning (DL) models and the XGBoost baseline were evaluated on the held-out test set to determine their predictive capabilities for resource requirement classification, as summarized in Table 2. While XGBoost—with label-encoded categorical features—provides a robust empirical benchmark with competitive accuracy and precision, the DL models demonstrate similar or slightly higher performance. Furthermore, the DL framework offers greater extensibility by incorporating additional features and capturing richer representations for future enhancements.
Model Metric XGBoost Deep Learning
Model 1 Accuracy 0.80 0.88
Macro F1 0.60 0.73
ROC-AUC 0.97 0.98
PR-AUC 0.92 0.96
Model 2 Accuracy 0.85 0.86
Macro F1 0.72 0.73
ROC-AUC 0.98 0.99
PR-AUC 0.95 0.95
Model 3 Accuracy 0.88 0.94
Macro F1 0.88 0.94
ROC-AUC 0.96 0.99
PR-AUC 0.96 0.98
Model 4 Accuracy 0.83 0.91
Macro F1 0.67 0.80
ROC-AUC 0.98 0.99
PR-AUC 0.93 0.97
Micro-average AUC for multiclass models.
Table 2: Summary of Performance Comparison of XGBoost and Deep Learning Models
@@
To further evaluate discriminative capability, we use the area under the receiver operating characteristic curve (ROC-AUC), with micro-averaged values provided in Table 2. In the multiclass context, this gives a global view of the model’s ability to separate all classes across threshold settings. Figure 5 shows the ROC curves for all models, visually demonstrating the quality of classification boundaries throughout the resource estimation pipeline. For binary classification tasks, including the I/O intensity model, the ROC curve provides insight into the model’s performance independent of class distribution [16].
Model 1 (RAM classification) attains 88% accuracy and macro F1 of 0.73, with a high ROC-AUC of 0.98. Notably, classwise results in Table 3 indicate the model achieves especially strong recall in bin1 (0.95) and bin4 (0.92). Model 2 (CPU time classification) matches this overall performance, with 86% accuracy, macro F1 of 0.73, and ROC-AUC of 0.99; the lowest per-class F1 scores (bin3: 0.52, bin4: 0.51) are due to class rarity but remain operationally useful.
Model 3, a binary classifier for I/O intensity, obtains 94% accuracy and macro F1 of 0.94, along with a near-perfect ROC-AUC of 0.99. Both “low” and “high” I/O intensity classes are identified robustly, as confirmed by the ROC curve in Figure 5c. For Model 4 (walltime classification), accuracy reaches 91% with macro F1 of 0.80 and ROC-AUC of 0.99, indicating uniformly high performance across the five prediction bins.
@@ 5 System Integration and Workflow Impact
We have developed a custom service-oriented architecture for integrating machine learning–driven resource prediction into the PanDA workload management system. The architecture is currently being evaluated in a dedicated test environment that closely replicates production conditions, enabling thorough validation with realistic scientific workloads. This test environment is essentially a reproduction of the production setup where real-time data used in production by the traditional Scout model is concurrently utilized by the machine learning pipeline for predictions. As depicted in Figure 6, while the traditional Scout model continuously operates in the production environment, the machine learning pipeline ingests the same live data within the test environment. This setup enables direct comparison and validation of the ML pipeline’s predictive performance and reliability against the benchmark traditional system under identical workload conditions, ensuring rigorous and practical evaluation prior to full-scale deployment. As shown in Figure 6, machine learning components are exposed as microservices within the workflow, deployed in containers to ensure consistent and reproducible execution. This approach facilitates automated resource prediction, precise scheduling, and streamlined workflow progression, with results from the test environment informing future production deployment.
@@ 5.1 Integrated Pipeline Evaluation for PanDA Brokerage
To quantify the effectiveness of our integrated solution, the PanDA brokerage system employs four predictive deep learning models that operate concurrently, providing real-time resource allocation estimates for incoming tasks. Assessing the overall system performance requires joint evaluation metrics that capture the collective contributions and interactions of these models within the broader workflow. Such an approach enables a holistic understanding of how the predictive ensemble drives improvements in scheduling accuracy, resource utilization, and workflow efficiency [14]. This integrated evaluation informs future refinements and will guide deployment in the full production environment.
@@ 5.2 Performance Benchmark and Workflow Impact
The integration of machine learning into the PanDA workflow system delivers transformative performance improvements. With sub-second inference latency, the ML pipeline achieves an orders-of-magnitude speedup over the traditional Scout-based approach, which required hours of task execution to estimate resource needs. Table 5 summarizes the key performance metrics:
@@ 6 Conclusion
In this study, we developed and rigorously evaluated machine learning models for resource requirement prediction in scientific workflows. Gradient-boosted decision trees (XGBoost) served as a strong empirical baseline, while deep neural networks with learned embeddings for categorical metadata achieved enhanced predictive accuracy and generalization. The use of adaptive embeddings and extensive regularization allowed the deep models to capture hidden relationships in workflow metadata, outperforming the baseline in both predictive stability and scalability. Comprehensive evaluation confirmed that the deep learning framework is well-suited for production integration, providing a reliable foundation for operational deployment within large-scale computing systems.
This work represents a significant advancement in intelligent resource management for large-scale scientific workflows. By providing accurate, early-stage predictions, our models help streamline resource allocation, reduce failures, and improve overall system throughput. Future work will focus on enhancing model adaptability across diverse workflow types and incorporating feedback mechanisms for continual model refinement.